OpenStudy (anonymous):

There were 170 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $490. How many Juniors bought tickets? Answer 90 70 80 100

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $490. How many Juniors bought. Tickets cost $1 and $2, or $5 per family at the door. ...... son is gluten-intolerant, as are many of her friends.

OpenStudy (anonymous):

so??? the answer is lol

OpenStudy (sasogeek):

the first question... i believe there's something missing... is all you wrote that there is to the question?

OpenStudy (lilg132):

490 = x + x 490 = 3.5x + 2x 490 = 3.5x + 2x 490 = 5.5x 89.09 = x 490 = 3.5(89.09) + 2(89.09) 490 = 311.815 + 178.18 490 = 312 + 178 89 people bought the senior tickets costing $3.5 totalling to $312 and 89 people bought the junior tickets costing $2 totalling to $178 I can see the closest answer to your given answers is 90 people so i would go with that dont forget to click good answer :)